IF A TRUCK LEFT TRAVELING NORTH AT 80 KM/H. AND ONE HOUR LATER ANOTHER TRUCK LEFT TRAVELING SOUTH AT 60 KM/H. HOW MANY HOURS HAD THE FIRST TRUCK TRAVELED WHEN THEY WERE 150 KM APART
MY EMAIL ADDRESS IS JMONAE96@HOTMAIL.COM

Answer by palanisamy(496)

A TRUCK LEFT TRAVELING NORTH AT 80 KM/H.
AND ONE HOUR LATER ANOTHER TRUCK LEFT TRAVELING SOUTH AT 60 KM/H.
After a time t hours, after the first truck had left,
The distance travelled by the first truck = 80t kms
The distance travelled by the second truck = 60(t-1) kms
Given the distance between them is 150 kms
80t+60(t-1) = 150
80t+60t-60 = 150
140t = 150+60
140t = 210
t = 210/140
t = 1.5 hours
t = 1 h 30 m